# On the Evaluation of Generative Models in Distributed Learning Tasks

### v3

@@ 4.1 Aggregate FID Scores in Distributed Learning
    For FID-all, if we define random $\widehat{X}$ with the average mean $\widehat{\bm{\mu}}=\sum_{i=1}^{k}\lambda_{i}\bm{\mu}_{i}$ and covariance matrix $\widehat{C}=\sum_{i=1}^{k}\lambda_{i}\bigl{(}C_{i}+\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigr{)}$ in the Inception-based semantic space, we have
@@
    For FID-avg, if we define a random vector $\widetilde{X}$ with the average mean $\widehat{\bm{\mu}}=\sum_{i=1}^{k}\lambda_{i}\bm{\mu}_{i}$ and covariance matrix $\widetilde{C}$ as the unique solution to $\widetilde{C}=\sum_{i=1}^{k}\lambda_{i}\bigl{(}\widetilde{C}^{1/2}C_{i}\widetilde{C}^{1/2}\bigr{)}^{1/2}$ in the Inception-based semantic space we will have
    $\mathrm{FID}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,$
    $=\,\mathrm{FID}\bigl{(}P_{\widetilde{X}},P_{G}\bigr{)}+\sum_{i=1}^{k}\lambda_{i}\mathrm{FID}\bigl{(}P_{\widetilde{X}},P_{X_{i}}\bigr{)}.$
    Thus, the FID-all as a function of $G$ is changing monotonically with $\mathrm{FID}\bigl{(}P_{\widetilde{X}},P_{G}\bigr{)}$ for defined $\widetilde{X}$.
@@
In Theorem 1, $\widetilde{C}$ follows from the Wasserstein barycenter of the Gaussian distributions and under the condition that $C_{1},\ldots,C_{k}$ commute, i.e. $C_{i}C_{j}=C_{j}C_{i}$ for every $i,j$, simplifies to
$\widetilde{C}=\Bigl{(}\sum_{i=1}^{k}\lambda_{i}C_{i}^{{1}/{2}}\Bigr{)}^{2}.$
Remark 3.
In Theorem 1, the optimal mean vectors for FID-all and FID-avg aggregations are the same. In contrast, the optimal covariance matrix of FID-avg denoted by $\widetilde{C}$ has no dependence on the choice of $\bm{\mu}_{i}$’s, while the optimal covariance matrix of FID-all $\widetilde{C}$ will be affected by the difference between $\bm{\mu}_{i}$’s due to the term $\sum_{i=1}^{k}\lambda_{i}(\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top})$. In general, Theorem 1 implies that the gap between FID-all and FID-avg can be written in the following form where $\mathrm{const.}$ remains constant under different $P_{G}$’s and $C_{G}$ is the embedded covariance matrix of $P_{G}$:
$\mathrm{FID}_{\mathrm{avg}}\bigl{(}P_{X_{1}},\ldots,P_{X_{k}};P_{G}\bigr{)}-\mathrm{FID}_{\mathrm{all}}\bigl{(}P_{X_{1}},\ldots,P_{X_{k}};P_{G}\bigr{)}\>$
= $=\>2\mathrm{Tr}\bigl{(}(C_{G}\widehat{C})^{1/2}-(C_{G}\widetilde{C})^{1/2}\bigr{)}\,+\,\mathrm{const.}$
As explained in the remarks, Theorem 1 shows that the optimal covariance matrices under FID-all and FID-avg could be significantly different in heterogeneous settings with different $\bm{\mu}_{i}$’s. Therefore, since the FID-all and FID-avg scores can be interpreted as the distance to covariance matrices $\widehat{C}$ and $\widetilde{C}$, respectively, the rankings suggested by the aggregations will be different if two generators’ covariances $C_{G}$ and $C_{G^{\prime}}$ have different ordering of distances to $\widehat{C}$ and $\widetilde{C}$.
4.2 KID-based Evaluation in Distributed Learning
After showing the possibility of inconsistent rankings by FID-all and FID-avg, we consider the KID score and analyze the consistency of KID-all and KID-avg aggregations. The following theorem proves that unlike the FID-case, the KID-all and KID-avg will result in a consistent ordering of the models and there is a monotonic relationship between the two aggregate scores.
Theorem 2.
$\mathrm{KID}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,$
$=\,\mathrm{KID}_{\mathrm{all}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}+\sum_{i=1}^{k}\lambda_{i}\,\mathrm{KID}\bigl{(}\widehat{P}_{X},P_{X_{i}}\bigr{)},$
Therefore, the above theorem shows the consistent rankings implied by the aggregate KID-avg and KID-all scores, as the difference between the scores remains constant while changing the model $P_{G}$.
@@ 5 Numerical Results
5.1 Evaluation on Synthetic Gaussian Mixture Datasets
As discussed in Remark 3, the optimal selection of the covariance matrix differs for the FID-all and FID-avg aggregate scores. To illustrate this distinction, we performed a toy experiment, revealing that FID-avg attains its minimum value when the generator’s variance closely approximates that of an individual client, whereas FID-all is minimized when the variance equals that of the aggregate distribution.
[…]
5.2 Evaluation on Real Image Datasets
We evaluated our theoretical results on standard image datasets. In our experiments, we simulated heterogeneous federated learning experiments consisting of non-i.i.d. data at different clients: for CIFAR-10 (Krizhevsky et al., 2009), we considered 10 clients, each owning 5000 samples exclusively from a single class of the image dataset. Therefore, every client’s dataset contains 5000 images having the same label. We put the experiment on CIFAR-100 and ImageNet-32 (Deng et al., 2009; Chrabaszcz et al., 2017) in Appendix.
Neural Net-based Generators. We trained a DDPM models using the standard training protocols outlined in (Ho et al., 2020) under the federated CIFAR-10 setting described above. To train the generative models in the federated learning setting, we employed the standard FedAvg approach (McMahan et al., 2017). Detailed information regarding the experiment setting can be found in the Appendix.
Figure 4: Illustration of randomly-generated samples from the variance-controlled generators. Images in each row are synthesized with the same truncation parameter.
@@ 5.2 Evaluation on Real Image Datasets
(c) Ranking
Figure 6: Evaluation on FFHQ with StyleGAN.
FID-based and KID-based Evaluation of Generative Models. We evaluated the generative models according to FID-all, FID-avg, KID-all, and KID-avg as defined in Section 4. In several cases, we observed that FID-all / FID-avg could assign inconsistent rankings to the generators. Specifically, we computed FID-all and FID-avg for the ideal ’airplane’-class-based generator and neural net-based DDPM generators under the distributed CIFAR10 setting. We present some examples generated from the two generators in Figure 3 and report their scores according to the four metrics. The results suggest that FID-avg assigns a considerably higher score to the ideal ’airplane’-based generator, whose images preserve perfect details but lack diversity in image categories. Conversely, FID-all assigns a relatively higher value to the DDPM model because its images possess greater diversity. On the other hand, we also observed that KID-avg and KID-all give consistent rankings. Both of them led to the evaluation that the ideal plane generator is slightly better than the DDPM generator. In our implementation of KID-scores, we utilized the standard implementation of KID measurement from data with a polynomial kernel, $k(\bf{x},\bf{y})=\left(\frac{1}{d}\bf{x}^{T}\bf{y}+1\right)^{3}$, where $d$ is the dimension of feature vector. We note that our theoretical finding on the evaluation consistency under KID-all and KID-avg applies to every kernel similarity function. We also test KID-scores with a Gaussian RBF kernel $k_{\sigma}^{\text{rbf}}(\bf{x},\bf{y})=\exp\left(-\frac{1}{2\sigma^{2}}\|\bf{x}-\bf{y}\|^{2}\right)$ as formulated in (Bińkowski et al., 2018), where we chose $\sigma=\sqrt{d}$ in the experiments. For images generated by diffusion model $\text{KID}^{\text{rbf}}$-all gives $4.277e^{-3}$ while $\text{KID}^{\text{rbf}}$-avg gives $4.295e^{-3}$. And for the airplane images in CIFAR10, $\text{KID}^{\text{rbf}}$-all gives $4.283e^{-3}$ while $\text{KID}^{\text{rbf}}$-avg gives $4.301e^{-3}$. The results indicate that for Gaussian RBF kernel $k_{\sigma}^{\text{rbf}}$, $\text{KID}^{\text{rbf}}$-all and $\text{KID}^{\text{rbf}}$-avg still gives consistent results. In this case, the $\text{KID}^{\text{rbf}}$-based evaluation suggests the images sampled from the diffusion model have higher quality than the set of airplane images in the CIFAR10 dataset.
Figure 7: Precision / Recall and Density / Coverage under variance-limited federated CIFAR10.
5.3 Evaluation of Well-Trained Generators in Distributed Learning Contexts
Here, we extend our numerical analysis to a federated learning scenario with realistic generators. We use the widely used StyleGAN2-ADA(Karras et al., 2020) generator, and download an FFHQ pre-trained model weight from the StyleGAN2-ADA’s GitHub repository. The generated images have the size $1024\times 1024$. We synthesized a sequence of variance-controlled generators by applying the standard truncation technique (Kynkäänniemi et al., 2019) on the random noise $\bf{z}$. The truncation parameter $\tau$ varies from $0.01$ to $1.0$. A greater truncation parameter $\tau$ leads to higher diversity in generated samples. We illustrate how the truncation parameter affects the samples in Figure 4. For each truncation $\tau$, we generate 5000 samples from the model for the evaluation. Within each client, we repeat the style vectors for the last eight layers in StyleGAN to get better visualization results, which will not affect the conclusion.
Furthermore, to simulate a distributed learning setting where each client only possesses a small set of variance-limited samples, we synthesized 100 clients. Each of the clients hold 100 images synthesized by the model with truncation $\tau=0.25$. The centre vector of each clients are various. The images within each client look similar, while images across clients look highly different as shown in Figure 5.
Similarly, we evaluate the models based on the two aggregations of FID and KID scores in this setting. The numerical results are shown in Figure 6a and Figure 6b. The FID-avg vs. FID-all plot leads to a U-shape curve, while the gap between KID-avg and KID-all remains constant for different generators. We further show the relation between the ranking based on FID and KID aggregate scores in Figure 6c
5.4 Precision/Recall and Density/Coverege Evaluations
In addition to FID and KID, we followed the definition in Equation 2 and Equation 3 and performed similar experiments to evaluate the consistency between the two aggregate scores for Precision/Recall (Kynkäänniemi et al., 2019) and Density/Coverage (Naeem et al., 2020). In the case of Precision/Recall, we utilized the official implementation with the number of clusters set to 5. We assessed precision and recall under the variance-limited CIFAR10 setting, with the results presented in Figure 7. The numerical scores indicate that in a heterogeneous data setting, the two aggregate precision scores may not consistently rank the generative models. On the other hand, based on the recall’s definition, it can be seen that Recall-all and Recall-avg will always take the same value, since the recall score reduces to an average over the generated data. Our numerical results are also consistent with this observation. For the Density/Coverage evaluations, the numerical results in Figure 7 suggest that while the density-based aggregate scores lead to more consistent rankings under data heterogeneity, both density-all/-avg may still provide inconsistent rankings. On the other hand, the coverage-based evaluations mimic the recall-based evaluation and consistently rank the models.
@@ 6 Conclusion
In this paper, we studied the evaluation of generative models in heterogeneous distributed learning problems where the clients have different distributions. We discussed the challenges of evaluating the overall performance of a trained generative model with only one score and showed the inconsistent rankings of sensible aggregations of standard FID scores in the network. On the other hand, we demonstrated that the same extensions of KID offer the same ranking of generative models. Our theoretical and experimental results indicate that KID-avg can be computed efficiently under the privacy constraints in distributed learning problems, while preserving the KID-all-based ranking of generative models. A possible future direction for our work is to extend the theoretical study to other evaluation criteria such as precision/recall and density/coverage scores. Also, understanding the behavior of the aggregate score using non-arithmetic averaging could be useful for evaluating deep generative models in federated learning contexts.
@@
We note that our numerical study focuses on the applications of generative models to image datasets and the empirical conclusions may not apply to other standard types of data including text and audio data. Regarding the work’s broader impact, we note that our analysis could be connected to the fairness evaluation of generative models in distributed learning contexts, as it suggests evaluation metrics for the assessment of diversity in the generated data. The study of fairness and diversity for generative models is required for a principled deployment of generative models in sensitive machine learning applications.
@@ A.1 Proof of Theorem 1
    In the above, $(a)$ follows from the Wasserstein-based definition of FID distance. $(b)$ comes from the well-known closed-form expression of the 2-Wasserstein distance between Gaussian distributions (Villani et al., 2009). $(c)$ is the result of applying the weighted barycenter of vector $\bm{\mu}_{1},\ldots,\bm{\mu}_{k}$ that can be seen to be $\widehat{\bm{\mu}}$ and the weighted barycenter of positive semi-definite covariance matrices $C_{1},\ldots,C_{k}$ that has been shown to be the unique matrix $\widehat{C}$ that solves the equation $\widetilde{C}=\sum_{i=1}^{k}\lambda_{i}\bigl{(}\widetilde{C}^{1/2}C_{i}\widetilde{C}^{1/2}\bigr{)}^{1/2}$ (Rüschendorf & Uckelmann, 2002; Puccetti et al., 2020). $(d)$ is the direct consequence of the Wasserstein-based definition of the FID distance and the closed-form expression of the 2-Wasserstein distance between Gaussians. Therefore, the proof is complete.

### v4

@@ 4.2 FID-based Evaluation in Distributed Settings
    For FID-all, if we define random $\widehat{X}$ with mean $\widehat{\bm{\mu}}=\sum_{i=1}^{k}\lambda_{i}\bm{\mu}_{i}$ and covariance matrix $\widehat{C}=\sum_{i=1}^{k}\lambda_{i}\bigl{(}C_{i}+\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top}\bigr{)}$, then
@@
    For FID-avg, if we define $\widetilde{X}$ with mean $\widehat{\bm{\mu}}=\sum_{i=1}^{k}\lambda_{i}\bm{\mu}_{i}$ and covariance matrix $\widetilde{C}$ as the unique solution to $\widetilde{C}=\sum_{i=1}^{k}\lambda_{i}\bigl{(}\widetilde{C}^{1/2}C_{i}\widetilde{C}^{1/2}\bigr{)}^{1/2}$, then
    $\mathrm{FID}_{\mathrm{avg}}\Bigl{(}P_{X_{1}},\ldots,P_{X_{k}}\,;\,P_{G}\Bigr{)}\,=\,\mathrm{FID}\bigl{(}P_{\widetilde{X}},P_{G}\bigr{)}+\sum_{i=1}^{k}\lambda_{i}\mathrm{FID}\bigl{(}P_{\widetilde{X}},P_{X_{i}}\bigr{)}.$
@@
In Theorem 2, $\widetilde{C}$ represents the covariance of the Wasserstein barycenter of the Gaussian distributions. If $C_{1},\ldots,C_{k}$ commute, i.e. $\forall i,j:\>C_{i}C_{j}=C_{j}C_{i}$, it simplifies to $\widetilde{C}=\bigl{(}\sum_{i=1}^{k}\lambda_{i}C_{i}^{{1}/{2}}\bigr{)}^{2}$
In the above theorem, we show the optimal mean vectors for FID-all and FID-avg aggregations are the same. In contrast, the optimal covariance matrix of FID-avg denoted by $\widetilde{C}$ has no dependence on the choice of $\bm{\mu}_{i}$’s, while the optimal covariance matrix of FID-all $\widetilde{C}$ will be affected by the difference between $\bm{\mu}_{i}$’s due to the term $\sum_{i=1}^{k}\lambda_{i}(\bm{\mu}_{i}\bm{\mu}_{i}^{\top}-\widehat{\bm{\mu}}\widehat{\bm{\mu}}^{\top})$. Therefore, Theorem 2 shows the distance-minimizing covariance matrices for FID-all and FID-avg are different in heterogeneous settings where $\bm{\mu}_{i}$’s are not identical, which shows the existence of models inconsistently ranked by FID-all and FID-avg.
As the FID-avg aggregation may not provide the FID-all’s ranking, a natural question is whether there exists another aggregation of the individual $\mathrm{FID}(P_{X_{i}},P_{G})$ scores which maintains the ranking of centralized FID-all evaluation. The following proposition shows that the answer to this question is no and it is possible that a group of clients assign identical individual FID scores to two generative models $P_{G}$ and $P_{G^{\prime}}$, while $P_{G}$ and $P_{G^{\prime}}$ have different FID-all scores.
Proposition 1.
In Theorem 2’s setting, suppose there exists $i,j$ that $\bm{\mu}_{i}\neq\bm{\mu}_{j}$. Also, assume the number of clients is less than the dimension of Inception embedding. Then, if we consider the FID-all minimizing $\widehat{G}$ with $\widehat{\bm{\mu}}$ and $\widehat{C}$ parameters, there exists a different generator $G^{\prime}$ that shares the same client-assigned FID scores $\mathrm{FID}(P_{X_{i}},P_{G^{\prime}})=\mathrm{FID}(P_{X_{i}},P_{\widehat{G}})$ for all $i$, while
[…]
As a result, two generative models may receive the same FID scores from a group of clients, while their FID-all scores are different according to the same set of clients. Therefore, neither averaging nor any other aggregation of clients’ FID scores can always preserve the ranking of the FID-all metric. Note that this statement does not apply to the KID score as shown in Theorem 1.
Table 1: The individual FID-scores of each clients and two aggregated FID scores.
Client 1 Client 2 Client 3 Client 4 Client 5 Client 6 FID-avg FID-all
Generator 1 281.76 198.54 212.00 225.21 129.14 123.59 195.04 100.49
Generator 2 282.44 199.24 212.36 226.21 129.34 122.61 195.37 190.93
Difference 0.68 0.70 0.36 1.00 0.20 0.98 0.33 90.44
Figure 2: Visual examples of generators and clients in Table 1.
@@ 5 Numerical Results
In this section, we test the theoretical results discussed in the text on the distributed ranking of generative models using the FID and KID metrics. We apply the truncation technique [16] to simulate a distributed evaluation setting with heterogeneous data distributions. Utilizing the simulated distributed scenario, we conducted numerical experiments to assess our theoretical results on FID and KID-based distributed evaluation. Also, we examined an implication of our results on detecting mode collapse in a distributed setting using the FID-avg and KID-avg scores.
5.1 Distributed vs. Centralized Evaluation using the FID metric
We earlier discussed that under identical FID evaluation scores assigned by the clients to two models, the model’s centralized FID score (FID-all) could still take different values. This subsection details experiments conducted on the FFHQ dataset [18] to demonstrate the possibility of such an outcome in real distributed settings.
(a) Generator 1 ($\tau=0.01$)
(b) Generator 2 ($\tau=0.2$)
(c) Generator 3 ($\tau=0.4$)
(d) Generator 4 ($\tau=0.6$)
(e) Generator 5 ($\tau=0.8$)
(f) Generator 6 ($\tau=1.0$)
Figure 3: Illustration of generators with different truncation factors. $\tau$ is the truncation parameter.
@@
Figure 4: Illustration of simulated clients with heterogeneous distributions via truncation technique.
Experiment Setup. We utilized the pre-trained StyleGAN2-ADA [18] generator on FFHQ [18] dataset provided in the official GitHub repository. The model’s generated images were of size $1024\times 1024$. Regarding the generators, by applying the truncation technique [16] and setting the truncation factor $\tau$ to vary over {0.01, 0.6}, we obtained two generators with highly different diversity scores; a higher truncation factor results in greater diversity. To simulate clients with highly heterogeneous distributions, we again applied the truncation technique and generated image samples centered around the randomly selected mean vectors. While the generated images for each simulated client looked similar, the samples of one client could be completely different from those of other simulated clients. We show randomly-selected samples from the generators and clients in Figure 2. Such a heterogeneous distributed scenario could occur in an actual mobile network, where the users’ collected photos have a higher inter-client variance than intra-client diversity. We calculated the individual FID score for each generator-client pair, averaging these to determine the FID-avg as in Equation 2. All the data in the users’ dataset were then used to calculate the FID-all metric for each generator as in Equation 1.
Experimental Results. As shown in Table 1, the experiment led to the identification of six clients with negligible differences in client-based FID scores across all clients (FID scores’ difference bounded by 1.0), suggesting that the two generators have similar FID-based scores from the perspective of each of the six clients. Therefore, both generators also had similar FID-avg scores (195.04 vs. 195.37). However, the FID-all scores (based on the six clients) led to significantly different scores for the two generators (100.49 vs. 190.93). The numerical result confirms our theoretical statement that similar client-based FID scores does not guarantee similar centralized FID-all scores.
@@
(c) Ranking(FIDs)
(d) Ranking(FID×$\times$×KID)
Figure 5: The results of evaluating generators in Figure 3 over clients in Figure 4.
Figure 6: Mode collapse Detection in the distributed evaluation setting. FID-avg failed to detect mode collapse, while KID-avg succeeded.
5.2 Comparison between Rankings implied by FID-based and KID-based Distributed Scores
Our theoretical analysis shows that the distributed KID-avg changes monotonically with the centralized KID-all score. To validate this claim, we conducted experiments on the simulated FFHQ data [18]. We also tested the hypothesis on several other datasets including Gaussian mixture clients, AFHQ [18], CIFAR [19], and ImageNet [20, 21] discussed in the Appendix.
Experiment Setting. Following a similar methodology as described in the previous subsection, we synthesized a series of diversity-controlled generators by applying the standard truncation technique to the random noise vector $\mathbf{z}$ [16]. We varied the truncation factor $\tau$ over $[0.01,1.0]$. The effect of changing the truncation factor on the generated samples’ diversity is illustrated in Figure 3. For every attempted $\tau$, we generated 5k samples. Additionally, to simulate a distributed setting with heterogeneous data distributions, we simulated 100 clients, each with 100 images synthesized with truncation factor $\tau=0.25$. We evaluated the generators in the distributed setting of the simulated clients using both the discussed FID and KID-based evaluation scores.
Numerical Results. As shown in Figure 5a and Figure 5b, FID-avg and KID-avg led to different rankings of the models. The plot of FID-avg versus FID-all led to a U-shaped curve, revealing inconsistent rankings of the models. On the other hand, the difference between KID-avg and KID-all remained constant for the simulated generators, as shown in Theorem 1. These findings are also visible in the comparative rankings based on FID and KID scores in Figures 5c and 5d. The results suggest that although KID scores and FID-all perfer generators with higher truncation factors and diversity, FID-avg preferred generated data with limited diversity matching the bounded diversity level at each client.
Knowing that the averaged scores are less costly for computation in distributed settings, KID-avg seems to offer a reliable proxy of KID-all and FID-all for a cost-efficient distributed ranking of generative models.
5.3 Mode Collapse Detection in Distributed Settings via KID-avg
During the training of generative adversarial networks (GANs), mode collapse may happen when the model fails to capture the diversity in the true distribution[22]. We analyzed the behavior of KID-avg and FID-avg for detecting mode collapse. In our experiments, the target generative model is evaluated by distributed FID-avg and KID-avg scores at fixed intervals during training, utilizing the same 100 clients described in Section 5.2. We simulated an experimental setting that led to mode collapse where we changed the training data at an epoch and begin training over a single data point. We report the evaluation outcomes and provide examples of generated images in Figure 6. We also show the centralized evaluation score FID-all and KID-all as a reference.
As demonstrated in Figure 6, the evaluated FID-avg scores could not help with detecting mode collapse, whereas FID-all, KID-all, and KID-avg all revealed a sudden deterioration in the scores. As FID-all and KID-all scores are more expensive for computation in a distributed setting, KID-avg could provide a less costly metric to detect mode collapse.
@@ 6 Conclusion
In this paper, we studied the evaluation of generative models in distributed settings where the clients can have heterogeneous data distributions. We discussed the score aggregation for widely-used FID and KID metrics to find the ranking of a group of generators according to the centralized FID-all and KID-all evaluation. We proved that averaging clients’ KID scores will address this task for the KID-based evaluation. On the other hand, we showed that no aggregations of clients’ FID scores can always recover the ranking of a centralized FID evaluation of generative models. A potential future direction is to characterize the computational and communication costs of distributed ranking of generative models based on the centralized FID-all score. In addition, extending the study to other evaluation metrics in the literature will be an interesting topic for future study.
@@
Our numerical experiments were limited to image-based generative models, and their extension to other domains such as video and text will be insightful on the application of KID-avg score for an inexpensive distributed evaluation of generative models.
@@ A.1 Proof of Theorem 1
    In the above, $(a)$ follows from the Wasserstein-based definition of FID distance. $(b)$ comes from the well-known closed-form expression of the 2-Wasserstein distance between Gaussian distributions [23]. $(c)$ is the result of applying the weighted barycenter of vector $\bm{\mu}_{1},\ldots,\bm{\mu}_{k}$ that can be seen to be $\widehat{\bm{\mu}}$ and the weighted barycenter of positive semi-definite covariance matrices $C_{1},\ldots,C_{k}$ that has been shown to be the unique matrix $\widehat{C}$ that solves the equation $\widetilde{C}=\sum_{i=1}^{k}\lambda_{i}\bigl{(}\widetilde{C}^{1/2}C_{i}\widetilde{C}^{1/2}\bigr{)}^{1/2}$ [24, 25]. $(d)$ is the direct consequence of the Wasserstein-based definition of the FID distance and the closed-form expression of the 2-Wasserstein distance between Gaussians. Therefore, the proof is complete.